# Vertex Fitting In Low-Material Budget Pixel Detectors

### v1

@@ III.3 Algorithm testing
    Hit parameters at the RS, $(u_{i},z_{i},\phi_{i},\lambda_{i},k_{i})_{MC}$, were obtained from MC trajectories.
@@
    $(u_{i},z_{i})_{MC}$ were smeared by using pixel resolution, i.e., by adding a random offset derived from a uniform distribution with mean and standard deviation $(0,\sigma_{l})$. The $(\phi_{i},\lambda_{i})$ angles were tilted according to the errors in equations 5 and 6 for values of the MS pooled from a Gaussian distribution with $\hat{\mu}=0$ and $\sigma=\theta_{\text{MS}}$. The error on $k_{i}$ was obtained by drawing from a Gaussian distribution with mean zero and a standard deviation $\sigma_{k}$.
@@
    From point 1 and 2, $\Xi_{\text{meas}}$ was derived together with the weight matrix $W$.
@@
    The initial parameters $\textbf{v}_{o}=x_{o},y_{o},z_{o}$ were obtained as the average coordinates of the tracks intersection points. Since two tracks can have up to two intersections, the one that is met first when back propagating the track parameters from the RS to the target is retained in the calculation of the average value. If two tracks do not intersect, the point of closest approach is used. The vectors $\textbf{t}_{o}$ were extracted at the point of closest approach of the trajectories to $\textbf{v}_{o}$ whilst $k_{o}$ was directly obtained from the track reconstruction carried out before the vertex fitting.
  5. 5.
@@
Figure 4: Pull distributions (a-f) of the track parameters at the vertex: ($x,y,z,\phi,\lambda,k)_{\text{v}}$ for test trajectories obtained by following the steps described in the text.
A precise determination of the fit errors depends on the correct characterization of $\Xi_{\text{meas}}$ and its propagation, e.g., WOLIN1993493 Lund_2009 . In the present study, the covariance matrix of the vertex parameters was calculated by propagating $\Xi_{\text{meas}}$ from the RS to the vertex point by using equations 18. A key test of the algorithm developed in this section consists in plotting the pull distributions of the track parameters. The pull of a variable $X$ with expected value $\mu_{X}$ and standard error $\sigma_{X}$ is:
@@
If the error and the residual in eq. 19 are well characterized, the pulls are normally distributed.
Figure 4 shows the normal distributions for the pulls of all the track parameters at the vertex $(x,y,z,\phi,\lambda,k)_{\text{v}}$ as calculated by the present vertex fitting algorithm.
IV Comaprative study of error sources
In section III.1, the explicit form of the covariance matrix was given by including the contribution of MS, pixel resolution and energy losses at RS: $\Xi_{\text{meas}}=\text{diag}\left\{\sigma_{u}^{2},\sigma_{z}^{2},\sigma_{\phi}^{2},\sigma_{\lambda}^{2},\sigma_{k}^{2}\right\}$. In this section, the relative weights of these errors on the determinations of the fit vertex parameters and their uncertainties are discussed. Three different inclusive scenarios have been considered: (A) MS only; (B) MS and pixel resolution are both included; (C) all sources of errors (MS, pixel resolution and energy losses) are considered. These three scenarios are summarized in table 1.
It must be noted that the kernel of $\Xi_{\text{meas}}$ grows larger going from scenario (A) to (C) along with the dimensionality of the problem, see for example (BILLOIR1992139, ). For instance, if the energy loss and pixel resolution at the RS are ignored, MS remains the only source of uncertainty against which all measurements are fixed but $(\phi_{i},\lambda_{i})_{\text{meas}}$, i.e., $\Xi_{\text{meas}}=\text{diag}\left\{\sigma_{\phi}^{2},\sigma_{\lambda}^{2}\right\}$. These 6 angles can be fitted with 3 vertex variables $(x,y,z)_{\text{v}}$ thus simplifying eq. 16, see e.g., schenk2013vertex .
Table 1: Scenrarios A,B,C in the comparative study of the vertex fitting.
@@ IV Comaprative study of error sources
Panels (a-e) in figure 5 show the the deviations between MC vertex parameters of simulated Mu3e decays and those obtained from vertex fitting in scenarios (A) and (B), respectively. From figure 5(a-c), it can be seen that the fit accuracy for the determinations of $(x,y,z)_{\text{v}}$ does not improve when the pixel resolution is included in covariance matrix. However, a significant improvement is found on the determinations of $(\phi_{i},\lambda_{i})_{\text{v}}$, as shown in figure 5(d,e). The results for scenario (C) are statistically the same as for scenario (B). The former being the only case in which the fit attempts to optimize the track parameter $k$, see table 1. As expected, having neglected the energy losses at the RS, track curvatures do not vary significantly throughout the vertex fitting, as it can be seen in figure 5(f).
[…]
The invariant mass of simulated Mu3e decays, calculated from the fit vertex parameters in scenarios (A) and (C), is shown in figure 7. As expected, no significant difference is seen in the two fit scenarios. In fact, the magnitude of the invariant mass is dominated by the muon rest mass over which the fit accuracy has little leverage.
Figure 6: Deviations between the fit and MC momentum coordinates $p_{x}$, $p_{y}$ and $p_{z}$ panels (a-c) and total momenutm $P$ in panel (d) for scenario (A) [full gray] and (C) [empty red], respectively. In (a,b,c), the legends show the mean and standard deviation obtained from a Gaussian fit whilst the legend in (d) shows the average and stadard deviation of the distributions.
Figure 7: Rreconstructed invariant mass for vertices with $\chi^{2}<15$ and total momentum $P<4$ MeV/c, scenario (A) [full gray] and (C) [empty red], respectively.
V Conclusion
In this paper, a simple least-squares method has been described which can be applied to reconstruct decay vertices in experiments equipped with pixel detectors. The relative weights of 1) MS, 2) pixel resolution and 3) energy losses to the final reconstruction accuracy has been investigated in the case study of the Mu3e low-material budget pixel detector. The exhaustive errors treatment of the present study goes beyond the MS-only approximation showing a significant improvement of the fit accuracy when the intrinsic pixel resolution is accounted for. This should encourage a rigorous treatment of the pixel resolution in the development of future reconstruction algorithms concerning precise particle physics measurements at low-energy.
VI Acknowledgements
I am grateful to the STFC grant for supporting this work. I wish to thank the members of the Mu3e Software and Analysis group for providing the simulation and track reconstruction software behind this study. I also want to thank Joel Goldstein, Niklaus Berger and Gavin Hesketh for all the detailed and useful discussions about this work. I also thank Naik Paras and Andre Schoning for their careful reading and useful comments.
VII Appendix 1: forward propagation of track parameters
The map $h(\textbf{v},\textbf{t},k)$ propagates a trajectory from the vertex to the hit at the RS. In this section, its analytical expression is derived by starting from the propagation of a track in the transverse plane and then along the beam direction.
@@ VII Appendix 1: forward propagation of track parameters
In this study, trajectories are helices with symmetry axis $\hat{z}$ and transverse radius $R_{\perp}$ which sign is given by the charge $q$ of the particle, see eq. 2. We write $R_{\perp}=R$cos($\lambda$) such as:
@@
$\ \;\;\sigma_{1}=m\,\left[-R^{2}\,m^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\,{\mathrm{sin}\left(\mathrm{\phi_{v}}\right)}^{2}+R^{2}\,m^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\right.\\ \left.-2\,R^{2}\,m\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)-R^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\,{\mathrm{cos}\left(\mathrm{\phi_{v}}\right)}^{2}\right.\\ \left.+R^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}+2\,R\,m^{2}\,\mathrm{x_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)+2\,R\,m\,\mathrm{x_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)\right.\\ \left.+2\,R\,m\,\mathrm{y_{o}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)-2\,R\,m\,\mathrm{y_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)\right.\\ \left.+2\,R\,\mathrm{y_{o}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)-2\,R\,\mathrm{y_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)-m^{2}\,{\mathrm{x_{v}}}^{2}\right.\\ \left.-2\,m\,\mathrm{x_{v}}\,\mathrm{y_{o}}+2\,m\,\mathrm{x_{v}}\,\mathrm{y_{v}}-{\mathrm{y_{o}}}^{2}+2\,\mathrm{y_{o}}\,\mathrm{y_{v}}-{\mathrm{y_{v}}}^{2}\right]^{1/2}\\ \mathrm{}\\ \;\;\sigma_{2}=R\,m^{2}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)\\ \mathrm{}\\ \;\;\sigma_{3}=R\,m\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)\,.\end{array}$ (25)
@@
The propagation of helical trajectories along the $\hat{z}$ axis is characterized by the following equations VALENTAN2009728 :

### v2

@@ 3.3 Algorithm testing
    Hit parameters at the RS, $(u^{i},z^{i},\phi^{i},\lambda^{i},k^{i})_{MC}$, were obtained from MC trajectories.
@@
    The parameters were smeared by adding a random offset according to the errors in $\Xi_{\text{meas}}$.
@@
    The initial parameters $\textbf{v}_{o}=x_{o},y_{o},z_{o}$ were obtained as the average coordinates of the tracks intersection points or points of closest approach. Since two tracks can have up to two intersections, the one that is encountered first during the propagation from the RS to the target is retained in the calculation of the average value. The vectors $\textbf{t}_{o}$ were extracted at the point of closest approach of the trajectories to $\textbf{v}_{o}$ whilst $k_{o}$ was directly obtained from the track reconstruction carried out before the vertex fitting.
@@
A key test of the algorithm consists in plotting the pull distributions of the track vertex parameters obtained from the fit. The pull of a variable $X$ with expected value $\mu_{X}$ and standard error $\sigma_{X}$ is:
@@
If the equations 16 and 18 are correct, the derived pulls should follow a normal distribution. Figure 4 displays the normal distributions for the pulls of all the track parameters at the vertex, as calculated in this test.
Figure 4: Pull distributions (a-f) of the track vertex parameters: ($x,y,z,\phi,\lambda,k)_{\text{v}}$, as calculated in the test procedure.
4 A comparative study of error sources
[…]
Figure 5: Blue diamonds, vertex resolution as a function of the MS for $\sigma_{l}=l/\sqrt{(}12)/1000$. Red circles,vertex resolution as a function of the $\sigma_{l}$ for MS $=0.0001^{\circ}$. The two arrows indicate the values of $\sigma_{MS}$ and $\sigma_{l}$ according to the Mu3e experimental design.
It is important to note that the kernel of $\Xi_{\text{meas}}$ grows larger going from scenario (A) to (C) along with the dimensionality of the problem, see for example [9]. For instance, if the energy loss and pixel resolution at the RS are ignored, MS remains the only source of uncertainty against which all measurements are fixed but $(\phi_{i},\lambda_{i})_{\text{meas}}$, i.e., $\Xi_{\text{meas}}=\text{diag}\left\{\sigma_{\phi}^{2},\sigma_{\lambda}^{2}\right\}$. These 6 angles can be fitted with 3 vertex variables $(x,y,z)_{\text{v}}$ thus simplifying eq. 16, see e.g., [23].
Panels (a-e) in figure 6 show the the deviations between MC vertex parameters of simulated Mu3e decays and those obtained from vertex fitting in scenarios (A) and (B), respectively. From figure 6(a-c), it can be seen that the fit accuracy for the determinations of $(x,y,z)_{\text{v}}$ does not improve when the pixel resolution is included in $\Xi_{\text{meas}}$. However, a significant improvement is found on the determinations of $(\phi_{i},\lambda_{i})_{\text{v}}$, as shown in figure 6(d,e). The results for scenario (C) are statistically the same as for scenario (B). The former being the only case in which the fit attempts to optimize the track parameter $k$, as detailed in table 1. As expected, having neglected energy losses at the RS, track curvatures do not vary significantly throughout the vertex fitting, as it can be seen in figure 6(f).
[…]
The reconstructed invariant mass of simulated Mu3e decays in scenarios (A) and (C) is shown in figure 8. As expected, no significant difference is seen in the two fit scenarios. In fact, the magnitude of the invariant mass is dominated by the muon rest mass over which the fit accuracy has little leverage.
Table 1: Scenarios A,B,C in the comparative study of the present vertex fitting algorithm.
@@ 4 A comparative study of error sources
Figure 7: Deviations, at the vertex, between the fit and MC momentum coordinates $p_{x}$, $p_{y}$ and $p_{z}$ panels (a-c) and total momentum $P$ in panel (d) for scenario (A) [full gray] and (C) [empty red], respectively. In (a,b,c), the legends show the mean and standard deviation obtained from a Gaussian fit whilst the legend in (d) shows the average and standard deviation of the distributions.
Figure 8: Reconstructed invariant mass of Mu3e decays, scenario (A) [full gray] and (C) [empty red], for $\chi^{2}\leq 15$ and $P\leq 4$ MeV/c.
5 Conclusion
In this work, a fitting algorithm appositely designed for the reconstruction of decay vertices in light pixel detectors was developed and described. The study focused on the Mu3e low-material budget pixel detector, exploring the relative impact of multiple scattering, pixel resolution and energy losses on the vertex resolution. The results demonstrate a significant improvement in the fit accuracy when accounting for the resolution of the pixel sensors. This should encourage a rigorous consideration of detector intrinsic spatial resolution in the development of future reconstruction algorithms concerning precise particle physics measurements at low energy.
6 Acknowledgements
This work was supported by the STFC Consolidated Grant PPR10542. I wish to thank the members of the Mu3e Software and Analysis group for providing the simulation and track reconstruction software used in this study. Special thanks go Joel Goldstein, Niklaus Berger and Gavin Hesketh for their detailed and insightful discussions about this work and Naik Paras and Andre Schöning for their careful reading and valuable comments.
Appendix A Propagation of track parameters
The analytical expression of $h(\textbf{v},\textbf{t},k)$ is derived by starting from the propagation of a track in the transverse plane and then along the beam direction.
@@ Appendix A Propagation of track parameters
In the transverse plane, trajectories are helices with symmetry axis $\hat{z}$ and transverse radius $R_{\perp}$ which sign is given by the charge $q$ of the particle, see eq. 2. We write $R_{\perp}=R$cos($\lambda$) such as:
@@
$\ \;\;\sigma_{1}=m\,\left[-R^{2}\,m^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\,{\mathrm{sin}\left(\mathrm{\phi_{v}}\right)}^{2}+R^{2}\,m^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\right.\\ \left.-2\,R^{2}\,m\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)-R^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}\,{\mathrm{cos}\left(\mathrm{\phi_{v}}\right)}^{2}\right.\\ \left.+R^{2}\,{\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)}^{2}+2\,R\,m^{2}\,\mathrm{x_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)\right.\\ \left.+2\,R\,m\,\mathrm{x_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)+2\,R\,m\,\mathrm{y_{o}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)\right.\\ \left.-2\,R\,m\,\mathrm{y_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)+2\,R\,\mathrm{y_{o}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)\right.\\ \left.-2\,R\,\mathrm{y_{v}}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)-m^{2}\,{\mathrm{x_{v}}}^{2}-2\,m\,\mathrm{x_{v}}\,\mathrm{y_{o}}\right.\\ \left.+2\,m\,\mathrm{x_{v}}\,\mathrm{y_{v}}-{\mathrm{y_{o}}}^{2}+2\,\mathrm{y_{o}}\,\mathrm{y_{v}}-{\mathrm{y_{v}}}^{2}\right]^{1/2}\\ \mathrm{}\\ \;\;\sigma_{2}=R\,m^{2}\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{cos}\left(\mathrm{\phi_{v}}\right)\\ \mathrm{}\\ \;\;\sigma_{3}=R\,m\,\mathrm{cos}\left(\mathrm{\lambda_{v}}\right)\,\mathrm{sin}\left(\mathrm{\phi_{v}}\right)\,.\end{array}$ (25)
@@
The propagation of helical trajectories along the $\hat{z}$ axis is characterized by the following equations [21]: